one of two supplementary angles exceeds 5 times the other angle by 12 degrees. find the measurement of the smaller angle.
Answer by macston(5194)

.
S=smaller angle L=larger angle=5S+12
.
S+L=180
S+5S+12=180
6S=168
S=28
ANSWER: The smaller angle is 28 degrees
.
CHECK:
L=5S+12
L=5(28)+12
L=140+12
L=152
The larger angle is 152 degrees.
.
L+S=180
152+28=180
180=180
.